OpenStudy (anonymous):

find the exact solution of cos 7pi/8

OpenStudy (anonymous):

-0.125

OpenStudy (anonymous):

:( this is what the answer shows in the exam review guide for trig \[-\sqrt{1/2-1/\sqrt(3)}\]

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

what would you like me to do :)

OpenStudy (anonymous):

-1/2-1/3 ??? I would try looking it up

OpenStudy (anonymous):

i did :( my math department at OSU sucks and makes their own problems

OpenStudy (anonymous):

i did send that shami98 a message though :) I hope he knows

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

hint: you need to use the double angle formula

OpenStudy (anonymous):

i do? O:

OpenStudy (anonymous):

the cos double angle or half angle? :)

OpenStudy (anonymous):

half angle works too

OpenStudy (anonymous):

also the identity cos(x)=-cos(pi-x) might help